OpenStudy (anonymous):

how can you get the integral of dy/(x^2 + 9)^2 using trigonometric substitution?

OpenStudy (anonymous):

x^2+a^2 --------> x=a tan t

OpenStudy (anonymous):

u have a tutorial for this water

OpenStudy (anonymous):

Actually @mukushla can get this integral by using trigonometric substitution. And he did that...

OpenStudy (anonymous):

Yes I have but I don't think that I have solved it using trigonometric identity there...

OpenStudy (anonymous):

Let me check..

OpenStudy (anonymous):

how are you able to integrate this equation if the variable of integration is y ? but the variable in equation is x?

OpenStudy (anonymous):

Yes I have done this there.. http://openstudy.com/study#/updates/4ff68236e4b01c7be8c976f1

OpenStudy (anonymous):

but it is dx? how about dy?

OpenStudy (anonymous):

For dy I think there is no need to evaluate this integral..

OpenStudy (anonymous):

The whole term is just a constant if dy is there..

OpenStudy (anonymous):

actually the given is this int. dy/(x^2 +9^2).. i just have difficulty because it is dy..

OpenStudy (anonymous):

You can take \(\frac{1}{x^2 + 9}\) out of the integral..

OpenStudy (anonymous):

Check your question if there is dy then in place of x^2 there will y^2 in the denominator..

OpenStudy (anonymous):

yes actually that's what i am also thinking about .. but the given is dy and the denominator is x^2 ..

OpenStudy (anonymous):

Your question simply evauates to : \[\frac{1}{(x^2 + 9)^2} \int\limits dy \rightarrow \frac{y}{(x^2 + 9)^2} + C\]

OpenStudy (anonymous):

so ur saying that i will assume that 1/ (x^2 + 9)^2 as a constant?

OpenStudy (anonymous):

by the way thank you :)

OpenStudy (anonymous):

you gave me a BIG help :)

OpenStudy (lgbasallote):

it's weird though why it asked for trig sub